### v2

@@ 1 Introduction
The weighted variant of the FOMC problem, known as the weighted first-order model counting (WFOMC) problem, additionally expects a specification of positive and negative weights for predicates in the language, which are then used to assign weights to models. In WFOMC, the task is to compute the sum of the weights of the given sentence’s models. [^1] Surprisingly, WFOMC is also essential for the unweighted FOMC problem since, for example, Skolemization for FOMC [dBMD14], which is used to eliminate existential quantifiers, relies on the usage of (negative) weights. WFOMC has applications in Statistical Relational Learning (SRL) [GT07]. For instance, Markov Logic Networks (MLNs) [RD06], one of the most popular SRL formalisms, model relational data by a distribution over possible worlds, where the computation of the partition function (used in probabilistic inference in MLN) is reducible to WFOMC [dB11].
@@
A key issue regarding WFOMC is its computational complexity. Generally, there is no hope of devising algorithms for WFOMC that would scale polynomially with the size of the first-order logic sentences under reasonable assumptions [BdBGS15]. This is true even for very simple fragments such as the one-variable fragment $\text{FO}^{1}$, which contains sentences with at most one logical variable. Therefore, most of the focus in this area has been on identifying fragments of first-order logic for which WFOMC can be computed in time polynomial in the size of the domain (but not necessarily in the size of the sentence). The term coined for this kind of tractability by [dB11] is domain liftability—it is an analog to data complexity [Var82] from database theory.
@@
The first positive results regarding the tractability of WFOMC came from the two seminal papers [dB11, dBMD14], which together established domain liftability of the two-variable fragment of first-order logic $\text{FO}^{2}$. This was quickly complemented by a hardness result [BdBGS15], showing that WFOMC for the three-variable fragment $\text{FO}^{3}$ is not domain-liftable (under plausible complexity-theoretic assumptions). However, this does not mean that the frontiers of tractability cannot be pushed beyond $\text{FO}^{2}$. For instance, the two-variable fragment with counting quantifiers $\text{C}^{2}$ [GOR97] (e.g., the above sentence of a $k$-regular graph) with cardinality constraints (i.e., the restrictions to the number of true ground atoms for relations), which is strictly more expressive than $\text{FO}^{2}$, was shown to be domain-liftable in [Kuz21]. Other works [dBD12, WPWK23, WPWK24] proved domain-liftability when attaching ground unary literals (as part of the input along with the domain size) to sentences from these fragments.
@@
One strategy for discovering new tractable fragments is to add extra axioms to a fragment known to be tractable, which may or may not be finitely expressible in first-order logic. This approach was first explored in [KL18] where it was shown that adding a single functionality axiom to $\text{FO}^{2}$ results in a domain-liftable fragment. [^2] This fragment is contained in $\text{C}^{2}$, which can encode an arbitrary number of functionality constraints. Subsequent works extended domain liftability to the fragments $\text{C}^{2}$ + Tree [vBK23] and $\text{C}^{2}$ + LinearOrder [TK23], which are obtained by adding to $\text{C}^{2}$ an axiom specifying that a distinguished relation $R$ should correspond to an undirected tree or to a linear order, respectively. Recently, the work [MBS23] added several new tractable classes: $\text{C}^{2}$ + DAG, $\text{C}^{2}$ + Connected, and $\text{C}^{2}$ + Forest, which allow restricting a distinguished binary relation to be a directed acyclic graph or to be a connected undirected graph or to be a forest. Except for the linear-order and the functionality axiom, these axioms are not finitely expressible in first-order logic.
@@
To add the axioms discussed above to $\text{C}^{2}$ while preserving domain-liftability, existing works usually rely on directly translating specialized techniques from enumerative combinatorics. For instance, Kirchhoff’s matrix-tree theorem was used for the tree axiom in [vBK23]. Similarly, ideas used in existing combinatorial proofs for counting connected graphs and acyclic graphs on $n$-vertices were adapted to the WFOMC setting in [MBS23] to implement the connectedness and acyclicity axioms. This raises a natural question:
@@ 1.1 Our Contribution
In this paper, we take a route inspired by graph polynomials such as the Tutte polynomial [Tut54] and directed chromatic polynomials [Sta70]. Graph polynomials are algebraic expressions that encode various combinatorial properties of graphs, allowing us to capture information such as connectivity, colorability, and spanning substructures. In this work, we consider $\text{C}^{2}$ possibly extended by cardinality constraints and unbounded ground unary literals (which we will call $\text{C}^{2}$ with cardinality constraints for simplicity). We introduce similar polynomials for first-order logic sentences, capturing certain combinatorial properties of their models, and show that they can be computed efficiently for $\text{C}^{2}$ with cardinality constraints, which requires introduction of new algorithmic techniques for computation of WFOMC. This, in turn, allows us to extend the domain-liftability to WFOMC for several new axioms (as well as for all of the old ones) and enables the efficient computation of graph polynomials for graphs that can be encoded by certain first-order fragments.
@@
$connected_{k}(R)$ $G(R)$ has $k$ connected components Theorem 3.6 new ($connected_{1}(R)$ by [MBS23])
@@
$tree(R)$ $G(R)$ is a tree Theorem 3.10 [vBK23]
[…]
@@
$AC(R)$ $G(R)$ is acyclic Theorem 4.9 [MBS23]
$DT(R,Root)$ $G(R)$ is a directed tree such that the root interprets the unary predicate $Root$ as true Theorem 4.10 [vBK23]
$DF(R)$ $G(R)$ is a directed graph such that each weakly connected component is a directed tree Theorem 4.10 [MBS23]
$LO(R)$ $G(R)$ represents a linear order Theorem 4.11 [TK23]
@@
Besides using the idea of graph polynomials to solve several counting tasks efficiently, conversely, the polynomials for WFOMC bring new results on graph polynomials. We show that the Tutte polynomial [Tut54], the strict and non-strict directed chromatic polynomials [Sta70] can be recovered from our polynomials for WFOMC. This allows us to show that these graph polynomials can be computed in time polynomial in the number of vertices for any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence (the parameters of cardinality constraints and the number of ground unary literals are allowed to depend on the number of vertices). This result answers in the positive way the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph (see its definition in Example 2.10) can be done in time polynomial in the number of vertices.
@@ 1.2 Related Work
We build on several lines of research. First, we build on the stream of results from lifted inference literature which established domain-liftability of several natural fragments of first-order logic [Poo03, dSBAR05, dB11, GD11, dBMD14, BdBGS15, KKVP16, Kuz21]. These works started with the seminal results showing domain-liftability of $\text{FO}^{2}$ [dB11, dBMD14] and continued by showing domain-liftability of larger fragments such as S2FO2 and S2RU [KKVP16], U1 [KL18] or $\text{C}^{2}$ [Kuz21]. Second, we build on extensions of these works that added axioms not expressible in these fragments or even in first-order logic such as the tree axiom [vBK23], forest axiom [MBS23], linear order axiom [TK23], acyclicity axiom [MBS23] and connectedness axiom [MBS23]. In this work, we generalize all of these results and add new axioms to this list, showing their domain-liftability. Finally, our work builds on tools from enumerative combinatorics [SS86] and on graph polynomials [Tut04]. The most well-known among graph polynomials is the Tutte polynomial [Tut54] but there exist also extensions of it [AB20], with which we do not work directly in this paper. Computing the Tutte polynomial is known to be intractable in general [JVW90], however, in this work we show that it can be computed in polynomial time for graphs that can be encoded by a fixed sentence from the first-order fragment $\text{C}^{2}$, possibly with cardinality constraints and ground unary literals, both of which may depend on the number of vertices of the graph (unlike the $\text{C}^{2}$ sentence which needs to stay fixed).
@@ 2.1 Mathematical Notations
Throughout this paper, we denote the set of natural numbers from 1 to $n$ as $[n]$. For a polynomial $f(x)$, the notation $[x^{i}]f$ represents the coefficient of the monomial $x^{i}$.
All vectors used in this paper are non-negative integer vectors. For a vector $\bm{c}$ of length $L$, $|\bm{c}|$ denotes the sum of its elements, i.e., $\sum_{i=1}^{L}c_{i}$. When considering two vectors $\bm{a}$ and $\bm{b}$ of the same length $L$, the operation $\bm{a}+\bm{b}$ denotes the vector $(a_{1}+b_{1},a_{2}+b_{2},\cdots,a_{L}+b_{L})$, and the subtraction is defined similarly. For a vector $\bm{a}$ and a non-negative integer $n$, $\binom{n}{\bm{a}}$ refers to $\binom{n}{a_{1}}\binom{n-a_{1}}{a_{2}}\cdots\binom{n-a_{1}-\cdots-a_{L-1}}{a_{L}}=\frac{n!}{a_{1}!a_{2}!\cdots a_{L}!}$.
@@ 2.3 Weighted First Order Model Counting
The weighted first-order model counting problem (WFOMC) takes the input consisting of a first-order sentence $\Psi$, a domain size $n\in\mathbb{N}$, and a pair of weighting functions $(w,\overline{w})$ that both map all predicates in $\Psi$ to real weights. Given a set $\mathcal{L}$ of ground literals whose predicates are in $\Psi$, the weight of $\mathcal{L}$ is defined as $W(\mathcal{L},w,\overline{w}):=\prod_{l\in\mathcal{L}_{T}}w(\mathsf{pred}\left(l\right))\cdot\prod_{l\in\mathcal{L}_{F}}\overline{w}(\mathsf{pred}\left(l\right)),$ where $\mathcal{L}_{T}$ (resp. $\mathcal{L}_{F}$) denotes the set of positive (resp. negative) ground literals in $\mathcal{L}$, and $\mathsf{pred}\left(l\right)$ maps a literal $l$ to its corresponding predicate name.
@@
Since these weightings are defined on the predicate level, all positive ground literals of the same predicate get the same weights, and so do all negative ground literals of the same predicate. For this reason, the notion of WFOMC defined here is also referred to as symmetric WFOMC.
@@
Consider $\Psi=\forall x\exists y\ R(x,y)$, and let $w(R)=1,\overline{w}(R)=1$. Then $\mathsf{WFOMC}(\Psi,n,w,\overline{w})=(2^{n}-1)^{n}$ because for each domain element $i\in[n]$, there are $12^{n}-$ 1 ways to choose the truth assignment of $R(i,1),R(i,2),\dots,R(i,n)$ that make $\exists y\ R(i,y)$ satisfied, and there are $n$ domain elements. For another example, consider the sentence $\Psi$ and the weighting functions $(w,\overline{w})$ in Example 2.1. One may check that $\mathsf{WFOMC}(\Psi,n,w,\overline{w})=(3^{n}+1)^{n}$.
@@ 2.4 Data Complexity of WFOMC
Definition 2.6 (Domain-liftability [dB11]).
@@
Domain-liftability of $\text{FO}^{2}$. Building on [dB11, dBMD14], in  [BdBGS15, Appendix C], it was shown that any sentence with at most two logical variables ($\text{FO}^{2}$) is domain-liftable by devising a polynomial-time algorithm for computing $\mathsf{WFOMC}(\Psi,n,w,\overline{w})$ for any $\text{FO}^{2}$ sentence $\Psi$ and any weighting $(w,\overline{w})$.
Here we briefly describe this algorithm, as some of its concepts are important in this paper. We mostly follow the notation from [vBK21]. Any sentence in $\text{FO}^{2}$ can be reduced to the prenex normal form with only universal quantifiers (i.e., $\forall x\forall y\ \psi(x,y)$, where $\psi(x,y)$ is a quantifier-free formula) by the normalization in [GKV97] and the technique of eliminating existential quantifiers in [dBMD14]. The reduction respects the WFOMC value, and hence in what follows, we assume that the input sentence is in the form $\forall x\forall y\ \psi(x,y)$. We first need the notion of a 1-type.
@@
Let $C_{1}(x),C_{2}(x),\cdots,C_{L}(x)$ be all the 1-types of $\Psi$. When we consider domain-liftability, the sentence $\Psi$ is fixed and therefore $L$ is a constant. We partition domain elements into 1-types by enumerating the configuration $\bm{c}=(c_{1},c_{2},\cdots,c_{L})$, a vector indicating the number of elements partitioned to each 1-type, since in the symmetric WFOMC only the configuration matters. Then we can rewrite $\Psi$ in the form of
$\Psi=\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}(\psi(x,y)\land\psi(y,x))\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi(x,y),$
@@
$\Psi=\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}\ \psi_{ij}(x,y)\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi_{i}(x,y).$
@@
$\mathsf{WFOMC}(\Psi,n,w,\overline{w})=\sum_{|\bm{c}|=n}\binom{n}{\bm{c}}\prod_{1\leq i\leq L}W(C_{i},w,\overline{w})^{c_{i}}\cdot s_{i}^{\binom{c_{i}}{2}}\prod_{1\leq i<j\leq L}r_{i,j}^{c_{i}c_{j}},$ (1)
@@
Domain-liftability of $\text{C}^{2}$. The above algorithm also serves as the basis of a WFOMC algorithm for $\text{C}^{2}$ (the $\text{FO}^{2}$ sentence with counting quantifiers) together with cardinality constraints and unbounded ground unary literals.
@@
For example, $|Eq|=n$ means that there are exactly $n$ positive ground literals of $Eq$ in a model, and the sentence $(\forall x\ Eq(x,x))\land|Eq|=n$ along with a domain of size $n$ encodes an identity relation $Eq$. Note that we allow cardinality constraints to depend on the domain size (such as in the definition of $Eq$ here where we used $|Eq|=n$).
@@
$\Psi_{B}=$ $\bigwedge_{V_{i},V_{j}\text{ are connected}}(Block_{i}(x)\land Block_{j}(y)\to E(x,y))$ (2)
$\land\bigwedge_{V_{i},V_{j}\text{ are not connected}}(Block_{i}(x)\land Block_{j}(y)\to\lnot E(x,y))$
$\land\bigwedge_{e=1}^{n}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right),$
@@
The presence of counting quantifiers, cardinality constraints, and ground unary literals does not affect the domain-liftability of $\text{FO}^{2}$. Using the techniques in [Kuz21] and [WPWK24, Appendix A], WFOMC of $\text{C}^{2}$ with cardinality constraints and unbounded ground unary literals can be reduced to WFOMC of $\text{FO}^{2}$ with symbolic weights [^6] The weighting functions $w,\overline{w}$ map predicates to symbolic variables. For more details, see [Kuz21, Section $3.1$]., and then solved by the algorithm for $\text{FO}^{2}$ in a straightforward manner.
@@
All the techniques mentioned above apply to the methods in this paper. Therefore, in the rest of this paper we only present the algorithms for $\text{FO}^{2}$ and state the results in terms of $\text{C}^{2}$ with cardinality constraints for simplicity.

### v3

@@ 1 Introduction
The weighted variant of the FOMC problem, known as the weighted first-order model counting (WFOMC) problem, additionally expects a specification of positive and negative weights for predicates in the language, which are then used to assign weights to models. In WFOMC, the task is to compute the sum of the weights of the given sentence’s models. [^1] Surprisingly, WFOMC is also essential for the unweighted FOMC problem since, for example, Skolemization for FOMC [WFOMC-FO2], which is used to eliminate existential quantifiers, relies on the usage of (negative) weights. WFOMC has applications in Statistical Relational Learning (SRL) [getoor2007introduction]. For instance, Markov Logic Networks (MLNs) [Richardson2006], one of the most popular SRL formalisms, model relational data by a distribution over possible worlds, where the computation of the partition function (used in probabilistic inference in MLN) is reducible to WFOMC [WFOMC-UFO2].
@@
A key issue regarding WFOMC is its computational complexity. Generally, there is no hope of devising algorithms for WFOMC that would scale polynomially with the size of the first-order logic sentences under reasonable assumptions [WFOMC-FO3]. This is true even for very simple fragments such as the one-variable fragment $\text{FO}^{1}$, which contains sentences with at most one logical variable. Therefore, most of the focus in this area has been on identifying fragments of first-order logic for which WFOMC can be computed in time polynomial in the size of the domain (but not necessarily in the size of the sentence). The term coined for this kind of tractability by [WFOMC-UFO2] is domain liftability—it is an analog to data complexity [datacomplexity] from database theory.
@@
The first positive results regarding the tractability of WFOMC came from the two seminal papers [WFOMC-UFO2, WFOMC-FO2], which together established domain liftability of the two-variable fragment of first-order logic $\text{FO}^{2}$. This was quickly complemented by a hardness result [WFOMC-FO3], showing that WFOMC for the three-variable fragment $\text{FO}^{3}$ is not domain-liftable (under plausible complexity-theoretic assumptions). However, this does not mean that the frontiers of tractability cannot be pushed beyond $\text{FO}^{2}$. For instance, the two-variable fragment with counting quantifiers $\text{C}^{2}$ [gradel1997two] (e.g., the above sentence of a $k$-regular graph) with cardinality constraints (i.e., the restrictions to the number of true ground atoms for relations), which is strictly more expressive than $\text{FO}^{2}$, was shown to be domain-liftable in [WFOMC-C2]. Other works [WFOMC-conditioning, WFOMS-FO2, WFOMS-FO2-new] proved domain-liftability when attaching ground unary literals (as part of the input along with the domain size) to sentences from these fragments.
@@
One strategy for discovering new tractable fragments is to add extra axioms to a fragment known to be tractable, which may or may not be finitely expressible in first-order logic. This approach was first explored in [DBLP:conf/lics/KuusistoL18] where it was shown that adding a single functionality axiom to $\text{FO}^{2}$ results in a domain-liftable fragment. [^2] This fragment is contained in $\text{C}^{2}$, which can encode an arbitrary number of functionality constraints. Subsequent works extended domain liftability to the fragments $\text{C}^{2}$ + Tree [WFOMC-tree-axioms] and $\text{C}^{2}$ + LinearOrder [WFOMC-linearorder-axiom], which are obtained by adding to $\text{C}^{2}$ an axiom specifying that a distinguished relation $R$ should correspond to an undirected tree or to a linear order, respectively. Recently, the work [WFOMC-axioms] added several new tractable classes: $\text{C}^{2}$ + DAG, $\text{C}^{2}$ + Connected, and $\text{C}^{2}$ + Forest, which allow restricting a distinguished binary relation to be a directed acyclic graph or to be a connected undirected graph or to be a forest. Except for the linear-order and the functionality axiom, these axioms are not finitely expressible in first-order logic.
@@
To add the axioms discussed above to $\text{C}^{2}$ while preserving domain-liftability, existing works usually rely on directly translating specialized techniques from enumerative combinatorics. For instance, Kirchhoff’s matrix-tree theorem was used for the tree axiom in [WFOMC-tree-axioms]. Similarly, ideas used in existing combinatorial proofs for counting connected graphs and acyclic graphs on $n$-vertices were adapted to the WFOMC setting in [WFOMC-axioms] to implement the connectedness and acyclicity axioms. This raises a natural question:
@@ 1.1 Our Contribution
In this paper, we take a route inspired by graph polynomials such as the Tutte polynomial [tuttepoly] and directed chromatic polynomials [chromatic-directed]. Graph polynomials are algebraic expressions that encode various combinatorial properties of graphs, allowing us to capture information such as connectivity, colorability, and spanning substructures. In this work, we consider $\text{C}^{2}$ possibly extended by cardinality constraints and ground unary literals (which we will call $\text{C}^{2}$ with cardinality constraints for simplicity). We introduce similar polynomials for first-order logic sentences, capturing certain combinatorial properties of their models, and show that they can be computed efficiently for $\text{C}^{2}$ with cardinality constraints, which requires the introduction of new algorithmic techniques for computation of WFOMC. This, in turn, allows us to extend the domain-liftability to WFOMC for several new axioms (as well as for all of the old ones) and enables the efficient computation of graph polynomials for graphs that can be encoded by certain first-order fragments.
@@
$connected_{k}(R)$ $G(R)$ has $k$ connected components Theorem 3.6 new ($connected_{1}(R)$ by [WFOMC-axioms])
@@
$tree(R)$ $G(R)$ is a tree Theorem 3.10 [WFOMC-tree-axioms]
$\mathit{forest}(R)$ $G(R)$ is a forest Theorem 3.11 [WFOMC-axioms]
@@
$AC(R)$ $G(R)$ is acyclic Theorem 4.9 [WFOMC-axioms]
$DT(R,Root)$ $G(R)$ is a directed tree such that the root interprets the unary predicate $Root$ as true Theorem 4.10 [WFOMC-tree-axioms]
$DF(R)$ $G(R)$ is a directed graph such that each weakly connected component is a directed tree Theorem 4.10 [WFOMC-axioms]
$LO(R)$ $G(R)$ represents a linear order Theorem 4.11 [WFOMC-linearorder-axiom]
@@
Besides using the idea of graph polynomials to solve several counting tasks efficiently, conversely, the polynomials for WFOMC bring new results on graph polynomials. We show that the Tutte polynomial [tuttepoly], the strict and non-strict directed chromatic polynomials [chromatic-directed] can be recovered from our polynomials for WFOMC. This allows us to show that these graph polynomials can be computed in time polynomial in the number of vertices for any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence (the parameters of cardinality constraints and the number of ground unary literals are allowed to depend on the number of vertices). This result answers in the positive way the question in [WFOMC-tree-axioms] whether calculating the Tutte polynomial of a block-structured graph (see its definition in Example 2.10) can be done in time polynomial in the number of vertices.
@@ 1.2 Related Work
We build on several lines of research. First, we build on the stream of results from lifted inference literature which established domain-liftability of several natural fragments of first-order logic [poole2003first, DBLP:conf/ijcai/BrazAR05, WFOMC-UFO2, DBLP:conf/uai/GogateD11a, WFOMC-FO2, WFOMC-FO3, DBLP:conf/nips/KazemiKBP16, WFOMC-C2]. These works started with the seminal results showing domain-liftability of $\text{FO}^{2}$ [WFOMC-UFO2, WFOMC-FO2] and continued by showing domain-liftability of larger fragments such as S2FO2 and S2RU [DBLP:conf/nips/KazemiKBP16], U1 [DBLP:conf/lics/KuusistoL18] or $\text{C}^{2}$ [WFOMC-C2]. Second, we build on extensions of these works that added axioms not expressible in these fragments or even in first-order logic such as the tree axiom [WFOMC-tree-axioms], forest axiom [WFOMC-axioms], linear order axiom [WFOMC-linearorder-axiom], acyclicity axiom [WFOMC-axioms] and connectedness axiom [WFOMC-axioms]. In this work, we generalize all of these results and add new axioms to this list, showing their domain-liftability. Finally, our work builds on tools from enumerative combinatorics [stanley1986enumerative] and on graph polynomials [tutte2004graph]. The most well-known among graph polynomials is the Tutte polynomial [tuttepoly] but there exist also extensions of it [tutte-directed], with which we do not work directly in this paper. Computing the Tutte polynomial is known to be intractable in general [tuttepoly-hardness], however, in this work we show that it can be computed in polynomial time for graphs that can be encoded by a fixed sentence from the first-order fragment $\text{C}^{2}$, possibly with cardinality constraints and ground unary literals, both of which may depend on the number of vertices of the graph (unlike the $\text{C}^{2}$ sentence which needs to stay fixed).
@@ 2.1 Mathematical Notations
Throughout this paper, we denote the set of natural numbers from 1 to $n$ as $[n]$. For a polynomial $f(x)$, the notation $x^{i}]f$ represents the coefficient of the monomial $x^{i}$.
All vectors used in this paper are non-negative integer vectors. For a vector $\bm{c}$ of length $p$, $|\bm{c}|$ denotes the sum of its elements, i.e., $\sum_{i=1}^{p}c_{i}$. When considering two vectors $\bm{a}$ and $\bm{b}$ of the same length $p$, the operation $\bm{a}+\bm{b}$ denotes the vector $(a_{1}+b_{1},a_{2}+b_{2},\cdots,a_{p}+b_{p})$, and the subtraction is defined similarly. For a vector $\bm{a}$ and a non-negative integer $n$, $\binom{n}{\bm{a}}$ refers to $\binom{n}{a_{1}}\binom{n-a_{1}}{a_{2}}\cdots\binom{n-a_{1}-\cdots-a_{p-1}}{a_{p}}=\frac{n!}{a_{1}!a_{2}!\cdots a_{p}!}$.
@@ 2.3 Weighted First Order Model Counting
The weighted first-order model counting problem (WFOMC) takes the input consisting of a first-order sentence $\Psi$, a domain size $n\in\mathbb{N}$, and a pair of weighting functions $(w,\overline{w})$ that both map all predicates in $\Psi$ to real weights. Given a set $L$ of ground literals whose predicates are in $\Psi$, the weight of $L$ is defined as $W(L,w,\overline{w}):=\prod_{l\in L_{T}}w(\mathsf{pred}\left(l\right))\cdot\prod_{l\in L_{F}}\overline{w}(\mathsf{pred}\left(l\right)),$ where $L_{T}$ (resp. $L_{F}$) denotes the set of positive (resp. negative) ground literals in $L$, and $\mathsf{pred}\left(l\right)$ maps a literal $l$ to its corresponding predicate name.
@@
Since these weightings are defined on the predicate level, all positive ground literals of the same predicate get the same weights, and so do all negative ground literals of the same predicate. For this reason, the notion of WFOMC defined here is also referred to as symmetric WFOMC [WFOMC-FO3].
@@
Consider $\Psi=\forall x\exists y\ R(x,y)$, and let $w(R)=1,\overline{w}(R)=1$. Then $\mathsf{WFOMC}(\Psi,n,w,\overline{w})=(2^{n}-1)^{n}$ because for each domain element $i\in[n]$, there are $2^{n}-1$ ways to choose the truth assignment of $R(i,1),R(i,2),\dots,R(i,n)$ that make $\exists y\ R(i,y)$ satisfied, and there are $n$ domain elements. For another example, consider the sentence $\Psi$ and the weighting functions $(w,\overline{w})$ in Example 2.1. One may check that $\mathsf{WFOMC}(\Psi,n,w,\overline{w})=(3^{n}+1)^{n}$.
@@ 2.4 Data Complexity of WFOMC
Definition 2.6 (Domain-liftability [WFOMC-UFO2]).
@@
Domain-liftability of $\text{FO}^{2}$. Building on [WFOMC-UFO2, WFOMC-FO2], in [WFOMC-FO3, Appendix C], it was shown that any sentence with at most two logical variables ($\text{FO}^{2}$) is domain-liftable by devising a polynomial-time algorithm for computing $\mathsf{WFOMC}(\Psi,n,w,\overline{w})$ for any $\text{FO}^{2}$ sentence $\Psi$ and any weighting $(w,\overline{w})$.
Here we briefly describe this algorithm, as some of its concepts are important in this paper. We mostly follow the notation from [WFOMC-FO2-faster]. Any sentence in $\text{FO}^{2}$ can be reduced to the prenex normal form with only universal quantifiers (i.e., $\forall x\forall y\ \psi(x,y)$, where $\psi(x,y)$ is a quantifier-free formula) by the normalization in [scott-normal-form] and the technique of eliminating existential quantifiers in [WFOMC-FO2]. The reduction respects the WFOMC value, and hence in what follows, we assume that the input sentence is in the form $\forall x\forall y\ \psi(x,y)$. We first need the notion of a 1-type.
@@
Let $C_{1}(x),C_{2}(x),\cdots,C_{p}(x)$ be all the 1-types of $\Psi$. When we consider domain-liftability, the sentence $\Psi$ is fixed and therefore $p$ is a constant. We partition domain elements into 1-types by enumerating the configuration $\bm{c}=(c_{1},c_{2},\cdots,c_{L})$, a vector indicating the number of elements partitioned to each 1-type, since in the symmetric WFOMC only the configuration matters. Then we can rewrite $\Psi$ in the form of
$\Psi=\bigwedge_{1\leq i<j\leq p}\forall x\in C_{i}\forall y\in C_{j}(\psi(x,y)\land\psi(y,x))\land\bigwedge_{1\leq i\leq p}\forall x\in C_{i}\forall y\in C_{i}\ \psi(x,y),$
@@
$\Psi=\bigwedge_{1\leq i<j\leq p}\forall x\in C_{i}\forall y\in C_{j}\ \psi_{ij}(x,y)\land\bigwedge_{1\leq i\leq p}\forall x\in C_{i}\forall y\in C_{i}\ \psi_{i}(x,y).$
@@
$\mathsf{WFOMC}(\Psi,n,w,\overline{w})=\sum_{|\bm{c}|=n}\binom{n}{\bm{c}}\prod_{1\leq i\leq p}W(C_{i},w,\overline{w})^{c_{i}}\cdot s_{i}^{\binom{c_{i}}{2}}\prod_{1\leq i<j\leq p}r_{i,j}^{c_{i}c_{j}},$ (1)
@@
Domain-liftability of $\text{C}^{2}$. The above algorithm also serves as the basis of a WFOMC algorithm for $\text{C}^{2}$ (the $\text{FO}^{2}$ sentence with counting quantifiers) together with cardinality constraints and ground unary literals.
@@
For example, $|Eq|=n$ means that there are exactly $n$ positive ground literals of $Eq$ in a model, and the sentence $(\forall x\ Eq(x,x))\land|Eq|=n$ along with a domain of size $n$ encodes an identity relation $Eq$. Note that we allow $k$ in Definition 2.9 to depend on the domain size (such as in the definition of $Eq$ here where we used $|Eq|=n$).
@@
$\Psi_{B}=$ $\bigwedge_{V_{i},V_{j}\text{ are connected}}\forall x\forall y(Block_{i}(x)\land Block_{j}(y)\to E(x,y))$ (2)
$\land\bigwedge_{V_{i},V_{j}\text{ are not connected}}\forall x\forall y(Block_{i}(x)\land Block_{j}(y)\to\lnot E(x,y))$
$\land\bigwedge_{e\in V}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right),$
@@
The presence of counting quantifiers, cardinality constraints, and ground unary literals does not affect the domain-liftability of $\text{FO}^{2}$. Using the techniques in [WFOMC-C2] and [WFOMS-FO2-new, Appendix A], WFOMC of $\text{C}^{2}$ with cardinality constraints and ground unary literals can be reduced to WFOMC of $\text{FO}^{2}$ with symbolic weights [^6] The weighting functions $w,\overline{w}$ map predicates to symbolic variables. For more details, see [WFOMC-C2, Section $3.1$]., and then solved by the algorithm for $\text{FO}^{2}$ in a straightforward manner.
@@
The techniques mentioned above apply to the methods in this paper. Therefore, in the rest of this paper we only present the algorithms for $\text{FO}^{2}$ and state the results in terms of $\text{C}^{2}$ with cardinality constraints for simplicity.